OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. 5 + yi = x + 3i

jimthompson5910 (jim_thompson5910):

Equate the real parts to get 5 = x and the imaginary parts to get yi = 3i

jimthompson5910 (jim_thompson5910):

From there, I'm sure you can see what x and y are

mathslover (mathslover):

I think @jim_thompson5910 meant to compare both the equations .. and the rest is right

mathslover (mathslover):

5 + yi = x + 3i |________| 5 + yi = x + 3i |_______| y = 3 and x = 5

mathslover (mathslover):

got it?

jimthompson5910 (jim_thompson5910):

that's a good way of looking at it (and displaying it) mathslover

mathslover (mathslover):

:) that waas what i was thinking thanks:)

OpenStudy (anonymous):

oh wow! Thanks to you both! I thought it was going to be more complicated!